A 60kg survivor of a shipwreck sits on a block of wood having volume of 1.6m³. Two birds of the block of the wood is under seawater (density = 1025 kg/m³ .find a the volume of diaplaced seawater by the block of wood, b. buoyant force on the block of wood and the survivor

1
Expert's answer
2022-04-11T12:44:35-0400

(a) the volume of displaced seawater by the block of wood

"V_d=\frac{m}{\rho_w}=\frac{60\:\rm kg/m^3}{1025\:\rm kg/m^3}=5.85*10^{-2}\:\rm m^3"

(b) the buoyant force on the block

"F_b=mg\\
=\rm 60\: kg*9.8\: N/kg=588\: N"